# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ I.2 Four-point functions
where $\left\langle\cdots\right\rangle_{\star D,\rho}$ and $\left\langle\cdots\right\rangle_{\star}$ are averages within the dynamic processes described by $W^{z}_{D,\rho}[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$ and $W^{x}[Q_{\star},0,R_{\star},\hat{R}_{\star}$], respectively. The yields neuronal and latent-variable single-site processes described by
$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (189)
$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (190)
@@
Isolating the upper-left element of the inverse of this matrix yields Eq. 51.
K.3 Sigmoidal parameterization of overlaps
To model structured left–right mode overlaps, we used the following sigmoidal form:
$\rho_{a}=\gamma_{\rho}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (199)
@@
While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective. This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix L. While these approaches yield equivalent results, they differ in their methodology: the path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables. Here we compute the four-point function using a two-site cavity approach, noting that the two-point function (which, if desired, could be computed using a single-site cavity method) is the same as that of an i.i.d. network with appropriate $g_{\text{eff}}$.
Leveraging Eq. 125, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al. [30] to both groups. This leads to separate, mutually referential cavity pictures for each group, mirroring the two single-site path integrals (Sec. IV). As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other. Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition (Eq. 25), the cavity method derives this function using its neuron-by-neuron definition (Eq. 24).
The cavity method we employ, schematized in Fig. 11, has two key features that distinguish it from simpler cavity calculations. First, it uses a two-site structure: we consider the simultaneous removal of two neurons, or two latent variables, from the network, rather than just one. This allows us to study cross-correlations between neurons. This approach was previously used in Clark et al. [30]. Second, it incorporates a bipartite structure: we treat both neurons and latent variables as dynamic objects, creating separate but mutually referential cavity pictures for each group. Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [96, 97]. Note that these two features are distinct: the two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
@@
as done in the path-integral approach. In the two-site cavity calculation, we keep track various intermediate quantities to order $1/\sqrt{N}$, leading to an expression for $\Psi^{\phi}(\bm{\tau})$ accurate to leading order (order one), which we identify as $\Psi^{\phi}_{\star}(\bm{\tau})$.
@@
$\delta z_{a}(t)=\int^{t}dt^{\prime}\sum_{b}S^{z}_{ab}(t_{1},t_{2})D_{b}\sum_{\mu}R_{\mu b}\phi_{\mu}(t^{\prime}),$ (202)
@@
$C^{\eta}_{00^{\prime}}(\omega)=\sum_{a,b}L_{0a}L_{0^{\prime}b}C^{z}_{ab}(\omega),$ (206)
@@
$C^{\phi}_{00^{\prime}}(\omega)=|S_{\star}^{\phi}(\omega)|^{2}C^{\eta}_{00^{\prime}}(\omega)+\frac{1}{\sqrt{N}}\left[(S_{\star}^{\phi}(\omega)F_{00^{\prime}}(\omega))^{*}+S_{\star}^{\phi}(\omega)F_{0^{\prime}0}(\omega)\right]C^{\phi}_{\star}(\omega).$ (207)
@@
where $\left\langle\cdot\right\rangle_{\bm{L},\bm{D},\bm{R}}$ denotes an average over $\bm{L}$, $\bm{D}$, and $\bm{R}$. To evaluate this, we need to square and disorder-average Eq. 207. This requires us to consider several two-frequency correlation functions, which we define as
@@
These have all been scaled to be order-one. We can evaluate these $\Gamma_{\cdots}(\bm{\omega})$ functions due to the independence of the couplings and dynamic variables in the expressions defining them, which is a consequence of the cavity construction. Of these, the only nonvanishing ones are
@@
$\delta\phi_{i}(t)=\int^{t}dt^{\prime}\sum_{j=1}^{N}S^{\phi}_{ij}(t_{1},t_{2})\sum_{\alpha}L_{j\alpha}z_{\alpha}(t^{\prime}),$ (215)
@@
$z_{\alpha}(t)=D_{\alpha}\left[\gamma_{\alpha}(t)+\frac{1}{\sqrt{N}}\sum_{\beta}\int^{t}dt^{\prime}G_{\alpha\beta}(t_{1},t_{2})z_{\beta}(t^{\prime})\right],$ (216)
@@
by squaring and disorder-averaging Eq. 220. This requires us to consider several two-frequency correlation functions, which we define as
@@
$S^{\phi}_{00^{\prime}}(\omega)$ $=\frac{1}{\sqrt{N}}(S^{\phi}_{\star}(\omega))^{2}F_{00^{\prime}}(\omega).$ (230)
Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 209-212 and 222225), we obtain
@@
We solve Eq. 232 and Eq. 234 simultaneously, yielding
@@
Now, we substitute Eq. 237 into Eq. 238. Using $\Psi^{\phi}_{\star}=\psi^{\phi}_{\star}+C^{\phi}_{12}$, we obtain
@@
where we used the previously defined quantities $g_{\text{eff}}^{2}=\alpha r_{2}$ and $\text{PR}^{D}=r_{2}^{2}/r_{4}$. Eq. 239 can be solved for $\Psi^{\phi}_{\star}$, yielding
@@
This result is identical to Eq. 38 derived using the path-integral approach, demonstrating the consistency between the two methods.
@@
To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$. For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (Sec. VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d. networks (see, e.g., [10, 98]). In summary:
@@
    Since $\eta(t)$ is Gaussian, so is $x(t)$ for this linear form of $T[\cdot]$, allowing us to write $C^{\phi}_{\star}(\tau)$ in terms of Gaussian integrals over $x(t)$ and $x(t+\tau)$, which have marginal variance $C^{\phi}_{\star}(0)$ and covariance $C^{\phi}_{\star}(\tau)$. Due to our choice of $\phi(x)$, this expression simplifies to
@@
    Squaring the single-site picture gives a second-order ODE $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$. Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
@@
    The rhs is integrated w.r.t. $C^{x}_{\star}(\tau)$ to provide an expression for $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$, which due to our choice of $\phi(x)$ simplifies to
@@
    Finally, Euler integration of the $C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq. 241.
We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$. When then use an FFT to obtain the frequency-space representation. The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
@@
with $\left\langle\phi^{\prime}\right\rangle_{\star}=1/\sqrt{1+(\pi/2)C^{x}_{\star}(0)}$. Various expressions for $\Psi^{a}_{\star}(\omega_{1},\omega_{2})$ were then computed in frequency space before transforming back to the time domain by a 2D inverse FFT.
For structured $\bm{L}$-$\bm{R}$ overlaps, the $\mathcal{O}(1)$ self-coupling in the single-site process makes $x(t)$ non-Gaussian, preventing the use of Gaussian integrals in the DMFT solution. We there obtain the solution via standard numerical methods, enforcing self-consistency among the $\eta^{x}(t)$ autocovariance function, $Q_{\star}(\tau)$; the self-coupling kernel, $R_{\star}(\tau)$; the response function, $S^{\phi}_{\star}(\tau)$; and the $\phi(t)$ autocovariance function, $C^{\phi}_{\star}(\tau)$. In summary:
@@
    $Q_{0}(\tau)$ and $R_{0}(\tau)$ are computed in Fourier space using the current values of $C^{\phi}_{0}(\tau)$ and $S^{\phi}_{0}(\tau)$ via Eq. 49, with the average taken over $(D,\rho)$.
@@
    Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq. 47a with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.
@@
    These time series $\phi(x_{s}(t))$ and $\eta^{x}_{s}(t)$ are used to compute empirical values $(C^{\phi}_{1}(\tau),S^{\phi}_{1}(\tau))$ for the autocovariance and response. The latter is computed by the Furutsu-Novikov theorem: $S^{\phi}(\omega)=C^{\phi\eta}(\omega)/C^{\eta}(\omega)$. The numerator is estimated using these time series, while $C^{\eta}(\omega)$ is simply $Q_{0}(\omega)$.
@@
We used a temporal discretization $\tau_{\text{max}}=120$, $d\tau=0.04$. With these quantities, the various components of Eq. 51 and the expression for $\Psi^{\phi}_{\star}$ are straightforward to compute. We took $\left\langle\cdot\right\rangle_{D,\rho}$ averages in this case by averaging over every other discrete value of $D_{a}$ and $\rho_{a}$.
@@
Each network was integrated via Runga-Kutta (RK4) with $dt=0.05$. The empirical lagged covariance matrix was computed by averaging over $N_{\text{loop}}$ individual covariance estimates, each of which is based on $T=2000$ discrete, evenly spaced time samples $t^{i_{\text{loop}}}_{k}$, with $t^{i_{\text{loop}}}_{k}-t^{i_{\text{loop}}}_{k-1}=1$:

### v3

@@ I.2 Four-point functions
where $\left\langle\cdots\right\rangle_{\star D,\rho}$ and $\left\langle\cdots\right\rangle_{\star}$ are averages within the dynamic processes described by $W^{z}_{D,\rho}[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$ and $W^{x}[Q_{\star},0,R_{\star},\hat{R}_{\star}$], respectively. This yields neuronal and latent-variable single-site processes described by
$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\circ\phi](t)$ (191)
$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\circ z](t),$ (192)
@@
Isolating the upper-left element of the inverse of this matrix yields Eq. (52).
K.3 Sigmoidal parametrization of overlaps
To model structured left-right mode overlaps, we used the following sigmoidal form:
$\rho_{a}=\gamma_{\rho}\left(1-\frac{2}{1+\exp[-\beta_{\rho}(a/N-u_{\text{flip}})]}\right),$ (201)
@@
While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective. This section provides an overview of the cavity analysis for the random-mode model. While these approaches yield equivalent results, they differ in their methodology: The path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables. Here, we compute the four-point function using a two-site cavity approach, noting that the two-point function (which, if desired, could be computed using a single-site cavity method) is the same as that of an i.i.d. network with appropriate $g_{\text{eff}}$.
Leveraging Eq. (127), we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al. [39] to both groups. This leads to separate, mutually referential cavity pictures for each group, mirroring the two single-site path integrals (Sec. IV). As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other. Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition [Eq. (25)], the cavity method derives this function using its neuron-by-neuron definition [Eq. (24)].
The cavity method we employ, schematized in Fig. 11, has two key features that distinguish it from simpler cavity calculations. First, it uses a two-site structure: We consider the simultaneous removal of two neurons, or two latent variables, from the network, rather than just one. This allows us to study cross-correlations between neurons. This approach was previously used in Clark et al. [39]. Second, it incorporates a bipartite structure: We treat both neurons and latent variables as dynamic objects, creating separate but mutually referential cavity pictures for each group. Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [103, 104, 69]. Note that these two features are distinct: The two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
@@
as done in the path-integral approach. In the two-site cavity calculation, we keep track of various intermediate quantities to order $1/\sqrt{N}$, leading to an expression for $\Psi^{\phi}(\bm{\tau})$ accurate to leading order (order one), which we identify as $\Psi^{\phi}_{\star}(\bm{\tau})$.
@@
$\delta z_{a}(t)=\int^{t}dt^{\prime}\sum_{b=1}^{M}S^{z}_{ab}(t_{1},t_{2})D_{b}\sum_{\mu\in\{0,0^{\prime}\}}R_{\mu b}\phi_{\mu}(t^{\prime}),$ (204)
@@
$C^{\eta}_{00^{\prime}}(\omega)=\sum_{a,b=1}^{M}L_{0a}L_{0^{\prime}b}C^{z}_{ab}(\omega),$ (208)
@@
$C^{\phi}_{00^{\prime}}(\omega)=|S_{\star}^{\phi}(\omega)|^{2}C^{\eta}_{00^{\prime}}(\omega)+\frac{1}{\sqrt{N}}\left[\left(S_{\star}^{\phi}(\omega)F_{00^{\prime}}(\omega)\right)^{*}+S_{\star}^{\phi}(\omega)F_{0^{\prime}0}(\omega)\right]C^{\phi}_{\star}(\omega).$ (209)
@@
where $\left\langle\cdot\right\rangle_{\bm{L},\bm{D},\bm{R}}$ denotes an average over $\bm{L}$, $\bm{D}$, and $\bm{R}$. To evaluate this, we need to square and disorder average Eq. (209). This requires us to consider several two-frequency correlation functions, which we define as
@@
These have all been scaled to be order one. We can evaluate these $\Gamma_{\cdots}(\bm{\omega})$ functions due to the independence of the couplings and dynamic variables in the expressions defining them, which is a consequence of the cavity construction. Of these, the only nonvanishing ones are
@@
$\delta\phi_{i}(t)=\int^{t}dt^{\prime}\sum_{j=1}^{N}S^{\phi}_{ij}(t_{1},t_{2})\sum_{\alpha\in\{\hat{0},\hat{0}^{\prime}\}}L_{j\alpha}z_{\alpha}(t^{\prime}),$ (217)
@@
$z_{\alpha}(t)=D_{\alpha}\left[\gamma_{\alpha}(t)+\frac{1}{\sqrt{N}}\sum_{\beta\in\{\hat{0},\hat{0}^{\prime}\}}\int^{t}dt^{\prime}G_{\alpha\beta}(t_{1},t_{2})z_{\beta}(t^{\prime})\right],$ (218)
@@
by squaring and disorder averaging Eq. (222). This requires us to consider several two-frequency correlation functions, which we define as
@@
$S^{\phi}_{00^{\prime}}(\omega)$ $=\frac{1}{\sqrt{N}}[S^{\phi}_{\star}(\omega)]^{2}F_{00^{\prime}}(\omega).$ (232)
Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions [Eqs. (211)–(214) and (224)–(227)], we obtain
@@
We solve Eqs. (234) and (236) simultaneously, yielding
@@
Now, we substitute Eq. (239) into Eq. (240). Using $\Psi^{\phi}_{\star}=\psi^{\phi}_{\star}+C^{\phi}_{12}$, we obtain
@@
where we used the previously defined quantities $g_{\text{eff}}^{2}=\alpha r_{2}$ and $\text{PR}^{D}=r_{2}^{2}/r_{4}$. Equation (241) can be solved for $\Psi^{\phi}_{\star}$, yielding
@@
This result is identical to Eq. (39) derived using the path-integral approach, demonstrating the consistency between the two methods.
@@
To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$. For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (Sec. VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d. networks (see, e.g., Refs. [10, 105]). In summary:
@@
    Since $\eta(t)$ is Gaussian, so is $x(t)$ for this linear form of $T[\cdot]$, allowing us to write $C^{\phi}_{\star}(\tau)$ in terms of Gaussian integrals over $x(t)$ and $x(t+\tau)$, which have marginal variance $C^{\phi}_{\star}(0)$ and covariance $C^{\phi}_{\star}(\tau)$. Because of our choice of $\phi(x)$, this expression simplifies to
@@
    Squaring the single-site picture gives a second-order ordinary differential equation (ODE) $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$. Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
@@
    The rhs is integrated with respect to $C^{x}_{\star}(\tau)$ to provide an expression for $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$, which due to our choice of $\phi(x)$ simplifies to
@@
    Finally, Euler integration of the $C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq. (243).
We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$. We then use an FFT to obtain the frequency-space representation. The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
@@
with $\left\langle\phi^{\prime}\right\rangle_{\star}=1/\sqrt{1+(\pi/2)C^{x}_{\star}(0)}$. Various expressions for $\Psi^{a}_{\star}(\omega_{1},\omega_{2})$ were then computed in frequency space before transforming back to the time domain by a two-dimensional inverse FFT.
For structured $\bm{L}$-$\bm{R}$ overlaps, the $\mathcal{O}(1)$ self-coupling in the single-site process makes $x(t)$ non-Gaussian, preventing the use of Gaussian integrals in the DMFT solution. We therefore obtain the solution via standard numerical methods, enforcing self-consistency among the $\eta^{x}(t)$ correlation function $Q_{\star}(\tau)$, the self-coupling kernel $R_{\star}(\tau)$, the response function $S^{\phi}_{\star}(\tau)$, and the $\phi(t)$ correlation function $C^{\phi}_{\star}(\tau)$. In summary:
@@
    $Q_{0}(\tau)$ and $R_{0}(\tau)$ are computed in Fourier space using the current values of $C^{\phi}_{0}(\tau)$ and $S^{\phi}_{0}(\tau)$ via Eq. (50), with the average taken over $(D,\rho)$.
@@
    Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq. (48a) with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi\bm{(}x_{s}(t)\bm{)}$.
@@
    These time series $\phi\bm{(}x_{s}(t)\bm{)}$ and $\eta^{x}_{s}(t)$ are used to compute empirical values $(C^{\phi}_{1}(\tau),S^{\phi}_{1}(\tau))$ for the autocovariance and response. The latter is computed by the Furutsu-Novikov theorem: $S^{\phi}(\omega)=C^{\phi\eta}(\omega)/C^{\eta}(\omega)$. The numerator is estimated using these time series, while $C^{\eta}(\omega)$ is simply $Q_{0}(\omega)$.
@@
We used a temporal discretization $\tau_{\text{max}}=120$, $d\tau=0.04$. With these quantities, the various components of Eq. (52) and the expression for $\Psi^{\phi}_{\star}$ are straightforward to compute. We took $\left\langle\cdot\right\rangle_{D,\rho}$ averages in this case by averaging over every other discrete value of $D_{a}$ and $\rho_{a}$.
@@
Each network was integrated via Runge-Kutta with $dt=0.05$. The empirical lagged covariance matrix was computed by averaging over $N_{\text{loop}}$ individual covariance estimates, each of which is based on $T=2000$ discrete, evenly spaced time samples $t^{i_{\text{loop}}}_{k}$, with $t^{i_{\text{loop}}}_{k}-t^{i_{\text{loop}}}_{k-1}=1$,